# KeyB2: Selecting Key Blocks is Also Important for Long Document Ranking with Large Language Models

### v1

@@ 4. Method: KeyB2
Based on the above conlusion, we show in this paper how to apply to LLMs the KeyB block selection approach proposed in (Li et al., 2023a) for long document IR. This answers our second research question regarding the design of block selection methods for LLM based IR rankers. The systems we propose, illustrated in Figure 5, are based on block selection mechanisms using fast models as standard IR models or PLMs, followed by LLM rankers based on Llama2 and Llama3, using a similar generation strategy as RankLLaMA.
This section introduces the long document segmentation approach, followed by three distinct block pre-ranking methods, including a novel one, and concludes with an aggregation of the top-scoring blocks.
@@ 4.2.1. KeyB2-BM25
BM25 is based on term matching mechanism. It improves TF-IDF approach which is also based on term matching. However, they adds some statistical information by calculating the inverse document frequency (IDF). Besides, compared with TF-IDF, BM25 introduces more parameters to modulate the impact of term frequency and takes into account the effects of document length and query length. Therefore, its performance is typically superior to that of TF-IDF in practical applications.
@@
$RSV(q,b)_{BM25}=\sum_{w\in q\cap b}IDF(w)\cdot\frac{tf_{w}^{b}}{k_{1}\cdot(1-b+b\cdot\frac{l_{b}}{l_{avg}})+tf_{w}^{b}},$
in which $tf_{w}^{b}$ corresponds to the number of occurrences of word $w$ in block $b$ and $IDF(w)$ to the inverse document frequency of word $w$. In this paper, we use the scikit-learn (Pedregosa et al., 2011) package to calculate the IDF dictionary with IDF smoothing. $IDF(w)$ is defined by:
@@
where $N$ and $df_{w}$ are added one with default scikit-learn smooth IDF setting. The added 1 after the fraction, makes sure terms with zero IDF don’t get suppressed entirely. Besides, $l_{b}$ is the length of block $b$, $l_{avg}$ the average length of the blocks in $d$, and $k_{1}$ and $b$ two hyperparameters.
@@ 4.2.2. KeyB2-cross
For each block $b_{i}$, to calculate its relevance with the query, we also can use a cross-encoder model, which should be efficient since the length of each block is limited. since the block number is relatively large, this model should be relatively fast. We use pre-trained language models (PLMs) like BERT (Devlin et al., 2019) as the cross-encoder, with a typical number of parameters ranging from dozens of millions to hundreds of millions, which is one to two orders of magnitude smaller than the one of LLMs. Denoting by $b_{i}$ the $i^{th}$ block and $b_{i}^{cls}$ the corresponding query-block representation, one has:
@@ 4.3. Aggregating the Local Relevance
5. Experiment
In this section, we evaluate the proposed approach on various datasets and compare it with baseline models. This section mainly aims at addressing RQ2 and RQ3 as introduced in Section 1.
@@ 5.1. Datasets
Since this paper focuses on document ranking, the datasets should be relatively long. In this paper, we use 3 datasets:
  • The first dataset is TREC 2019 DL document ranking task (Craswell et al., 2020). The Deep Learning Track, introduced for TREC 2019, aims to study large data ad hoc ranking. It is the first track large human-labeled training datasets. It includes two tasks: document ranking and passage ranking. In this paper, we focus on the document ranking task.
  • The second dataset is Robust04444https://trec.nist.gov/data/robust/04.guidelines.html, whose sources include Financial Times, Federal Register 94, FBIS, disk 5 and LA Times.
  • The last dataset MLDR Multilingual Long-Document Retrieval dataset (Chen et al., 2024), and we use the MLDR-zh (Chinese) subset. The source of MLDR-zh are Wikipedia and Wudao (Yuan et al., 2021).
@@ 5.2. Baselines
To evaluate the performance of KeyB2 models, we compare them against several baseline approaches across multiple datasets:
  • On TREC 2019 DL: Following (Li et al., 2023a), the baselines used for this dataset include traditional methods like BM25, neural IR models such as CO-PACRR (Hui et al., 2018), TK (Hofstätter et al., 2020), and TKL (Hofstätter et al., 2020), as well as transformer-based models like RoBERTa (FirstP and MaxP) (Liu et al., 2019) and PARADE (Li et al., 2020). Additionally, we compare with sparse attention models, including Sparse-Transformer (Child et al., 2019), Longformer-QA (Beltagy et al., 2020), Transformer-XH (Zhao et al., 2020), and QDS-Transformer (Jiang et al., 2020). Previous key block selection models like IDCM (Hofstätter et al., 2021) and KeyB (Li et al., 2023a) are also included. The LLM-based RankLLaMA (Ma et al., 2024) model serves as another important baseline.
  • On Robust04: The baselines for this dataset consist of BM25, as well as BERT-based KeyB models (using BM25 and BinB configurations). We also compare with the RankLLaMA model.
  • On MLDR-zh: the baselines include traditional IR methods such as BM25, along with KeyB models using BM25 and BinB configurations. For this dataset, the KeyB models utilize a Chinese version of BERT and a BM25 implementation adapted for Chinese text. The RankLLaMA model is also used as a baseline for multilingual retrieval.
@@ 5.3. Experimental Design and Implementation Details
Experimental Design
To evaluate KeyB2, we first fine-tuned the models on the TREC 2019 DL document ranking dataset, comparing them with various baselines. For a fair comparison with RankLLaMA, which uses the Llama2555https://huggingface.co/meta-llama/Llama-2-7b-hf model with 7B size, we initially experimented with KeyB2 using the same Llama2, referred to as KeyB2(Llama2). To explore potential improvements, we also tested KeyB2 with Llama3666https://huggingface.co/meta-llama/Meta-Llama-3-8B model with 8B size, named KeyB2(Llama3). The KeyB2 approach includes three key block selection methods: BM25, cross-encoder, and bi-encoder. We fine-tuned KeyB2(Llama2) and KeyB2(Llama3) for each method, resulting in configurations such as KeyB2(Llama2)BM25, KeyB(Llama2)cross and KeyB(Llama2)bi.
Next, we evaluated the zero-shot performance by applying each KeyB2 model, trained on the TREC 2019 DL dataset, to the Robust04 dataset without additional fine-tuning. Finally, we conducted other fine-tuning experiments, where both KeyB2 and baseline models were trained from scratch on the downstream datasets Robust04 and MLDR-zh. The tasks of all models excluding BM25 for the datasets are to rerank documents in the test sets, while BM25 shows the retrieval performances (Yang et al., 2018).
Implementation Details
We implemented KeyB2 and RankLLaMA using PyTorch 2.4.0 and CUDA 11.8, with training conducted on an NVIDIA A100 GPU. The main configurations and training details are as follows:
@@
    • The maximum query token number is 32. For KeyB2, the document token number is limited to 480 after block selection, resulting in a total input size of 512 tokens, matching the KeyB (Li et al., 2023a) limit.
    • For RankLLaMA, which processes documents directly, the maximum token number is set to 1200 to fit within A100 GPU.
@@
    Training Configurations:
    • For KeyB2 models, the batch size is set to 2. To efficiently fine-tune large language models (LLMs), we employ the LoRA technique (Hu et al., 2021), which freezes LLM weights and introduces trainable matrices to reduce the number of trainable parameters. LoRA parameters were set to $r=32$ and $\alpha=64$.
    • The learning rate is $5\times 10^{-5}$, with a gradient accumulation step of 8. For RankLLaMA, the batch size is 1 and other settings are the same with KeyB2. AdamW optimizer is used with WarmupDecayLR scheduler using DeepSpeed (Rasley et al., 2020), with FP16 mixed precision.
    • For KeyB models, following the code from (Li et al., 2023a) with the learning rate $5\times 10^{-5}$.
  3. (3)
    Model Variants for Different Languages:
    • KeyB2 uses small pre-trained models for cross-encoders or bi-encoders depending on the dataset language, as detailed in Table 1.
@@
    • For English datasets, BM25 is implemented using the Anserini tool (Yang et al., 2018) with default settings ($k_{1}=0.9$, $b=0.4$). For KeyB2-BM25, we computed the IDF dictionary using scikit-learn on the TREC 2019 DL dataset, with parameters matching Anserini’s defaults.
    • For the Chinese MLDR-zh dataset, the BM25 baseline is also implemented with Anserini, using default settings. For KeyB2-BM25, as mentioned in Section 4.2.1, the sentences are separated into Chinese words. Other settings are the same to English datasets.
  5. (5)
    Training Data:
    • TREC 2019 DL: 200K triplets of a query, a positive document, and a negative document.
    • Robust04: 5-fold cross-validation with 40K triplets per fold, and fold separation follows (Li et al., 2023a).
    • MLDR-zh: 10,000 training queries, each with a positive and negative document.
  6. (6)
    Fine-Tuning Approach:
    • Fine-tuning uses pairwise hinge loss with a learning rate of $5\times 10^{-5}$.
    • Models are initialized from official Llama checkpoints.
@@
    • On TREC 2019 DL, following previous works (Li et al., 2020, 2023a), NDCG@10 and MAP are used.
    • The evaluation metrics include precision at 1, 5, 10, and 20 (P@k), as well as NDCG@k, NDCG and MAP (Li et al., 2023a).
    • On MLDR-zh, in the qrel file of MLDR Chinese subset, each query corresponds to one positive label, and in the test file, each query consists 8 documents for ranking. So for the P@k metric, we report the P@1, indicating the percentage of the relevant document is ranked in the first position by the approaches. Besides, MAP is reported. For NDCG@k, we report values at NDCG@1, NDCG@5, and NDCG@8.
@@
TREC 2019 DL & Robust04 MLDR-zh
@@
5.4. Fine-tuning on TREC DL19 Document Ranking Task
Table 2. Experiment on TREC 2019 DL and comparison with sparse attention models and previous select blocks models. Baseline results before the LLM and KeyB2 parts are from (Li et al., 2023a). Best results are in bold.
TREC Deep Learning Track Document Ranking
@@ 5.4. Fine-tuning on TREC DL19 Document Ranking Task
RankLLaMA-original paper (Ma et al., 2024) 0.677 -
RankLLaMA-trainByUs1200 0.707 0.291
@@
KeyB2(Llama2)BM25 0.713 0.293
KeyB2(Llama2)cross 0.730 0.302
KeyB2(Llama2)bi 0.710 0.293
KeyB2(Llama3)BM25 0.718 0.293
KeyB2(Llama3)cross 0.731 0.298
KeyB2(Llama3)bi 0.711 0.296
The TREC DL19 document ranking task is representative for long document IR experiments as it includes a large amount of training data suitable for effective model training. The results demonstrate that the KeyB2 models outperform multiple baseline approaches on the TREC 2019 DL document ranking task, highlighting the effectiveness of the proposed approach. KeyB2 models, particularly those using Llama2 and Llama3 with cross-encoder configurations, consistently achieve higher NDCG@10 and MAP scores than baseline methods, sparse attention models, and prior key block selection techniques.
Performance Highlights:
  • KeyB2 vs. Traditional and Neural IR Baselines: The KeyB2 models achieve largely gains over traditional methods like BM25 and neural IR models such as CO-PACRR and PARADE. For example, KeyB2(Llama3)cross improves NDCG@10 by more than 30% over BM25 and by approximately 12% over PARADE. This indicates that the block selection strategy in KeyB2 effectively enhances the ranking of long documents by focusing on the most relevant content.
  • Comparison with Sparse Attention Models: KeyB2 outperforms sparse attention models such as Longformer-QA and QDS-Transformer, with NDCG@10 improvements of around 10%. This suggests that selecting key blocks is more effective than applying sparse attention mechanisms to the entire document.
  • Advantage over Previous Key Block Selection Methods: Compared to earlier key block selection models like KeyB(BERT), KeyB2(Llama3)cross increases the result in NDCG@10 from the best 0.707 to 0.731, and KeyB2(Llama3)BM25 increases the NDCG@10 from 0.678 to 0.718. The results also largely outperform KeyB(PARADE5). These demonstrate the benefit of integrating LLMs with the block selection approach.
  • KeyB2 vs. RankLLaMA: With the same Llama model, The KeyB2(Llama2)cross configuration improves the score in NDCG@10 from 0.677 (reported in (Ma et al., 2024)) to 0.730, or 0.707 (trained by us using pairwise Hinge loss) to 0.730. This indicating that focusing on selected blocks leads to better ranking accuracy than processing whole documents directly.

### v2

@@ 4.1. Recap of KeyB Framework and Differences in KeyB2
Overall, KeyB2 preserves the core idea of KeyB while extending the framework to the LLM era with stronger backbones, richer selector choices, and deeper empirical analysis. The systems we propose, illustrated in Figure 7, are based on block selection mechanisms using fast models as standard IR models or PLMs, followed by LLMs and trained as rankers. Next we introduce the document segmentation approach, followed by three distinct block pre-ranking methods, including a novel one, and concludes with an aggregation of the top-scoring blocks.
@@ 4.3.1. KeyB2-BM25
BM25 is a term-matching retrieval model that extends the classic TF-IDF formulation. While both methods down-weight ubiquitous terms through inverse document frequency, BM25 further (i) applies a sub-linear saturation function to term frequency, and (ii) introduces a length-normalization factor that penalizes disproportionately long documents. These two mechanisms are controlled by tunable hyper-parameters $k_{1}$ and $b$, enabling BM25 to balance term frequency impact and document length effects more effectively than TF-IDF, and thus to deliver stronger empirical performance in long-document ranking tasks.
@@
$RSV(q,blk)_{BM25}=\sum_{w\in q\cap blk}IDF(w)\cdot\frac{tf_{w}^{blk}}{k_{1}\cdot(1-b+b\cdot\frac{l_{blk}}{l_{avg}})+tf_{w}^{blk}},$
in which $tf_{w}^{blk}$ corresponds to the number of occurrences of word $w$ in the block $blk$ and $IDF(w)$ to the inverse document frequency of word $w$. In this paper, we use the scikit-learn (Pedregosa et al., 2011) package to calculate the IDF dictionary with IDF smoothing. $IDF(w)$ is defined by:
@@
where $N$ and $df_{w}$ are added one with default scikit-learn smooth IDF setting. The added 1 after the fraction, makes sure terms with zero IDF don’t get suppressed entirely. Besides, $l_{blk}$ is the length of block, $l_{avg}$ the average length of the blocks in $d$, and $k_{1}$ and $b$ two hyperparameters of BM25.
@@ 4.3.2. KeyB2-cross
For each block denoted as $b_{i}$, to calculate its relevance with the query, we also can use a cross-encoder model, which should be efficient since the length of each block is limited. Since the block number is relatively large, this model should be relatively fast. We use pre-trained language models (PLMs) like BERT (Devlin et al., 2019) as the cross-encoder, with a typical number of parameters ranging from dozens of millions to hundreds of millions, which is one to two orders of magnitude smaller than the one of LLMs. Denoting by $b_{i}$ the $i^{th}$ block and $b_{i}^{cls}$ the corresponding query-block representation, one has:
@@ 4.4. Aggregating the Local Relevance
5. Experimental Settings
From this section, we evaluate the proposed approach on various datasets and compare it with baseline models, mainly aiming at addressing RQ2, RQ3 and RQ4 as introduced in Section 1.
@@ 5.1. Datasets
Since this paper focuses on document ranking, the datasets should be relatively long. In this paper, we use 4 datasets:
  • TREC 2019 DL document reranking task (Craswell et al., 2020) is the first dataset. The Deep Learning Track, introduced for TREC 2019, aims to study large data ad hoc ranking. It is the first track large human-labeled training datasets. It includes two tasks: document ranking and passage ranking. In this paper, we focus on the document reranking task.
  • TREC 2023 DL document reranking task is further experimented on different models, whose corpus is based on MS MARCO v2, different with DL19’s.
  • Robust04666https://trec.nist.gov/data/robust/04.guidelines.html is the third dataset, whose sources include Financial Times, Federal Register 94, FBIS, disk 5 and LA Times.
  • MLDR Multilingual Long-Document Retrieval dataset (Chen et al., 2024) is the last dataset, and we use the MLDR-zh (Chinese) subset. The sources of MLDR-zh are Wikipedia and Wudao (Yuan et al., 2021).
Table 2 summarizes key statistics of the datasets, including document genres, corpus sizes, test query counts, and average document lengths (using Llama2 tokenizer). Among them, MLDR-zh features the longest documents, while Robust04 primarily contains shorter newswire articles. These datasets offer a diverse experimental setting for evaluating KeyB2.
Table 2. Datasets statistics
Dataset genre #documents #test query Avg. Num. of Tokens
TREC DL19 doc (MS MARCO) Web Documents 3,213,835 49 1958
TREC DL23 doc (MS MARCO v2) Web Documents 11,959,635 82 3782
Robust04 News Articles 528,030 250 741
MLDR-zh Wikipedia + Wudao 200,000 800 12899
@@ 5.2. Baselines
We compare KeyB2 against several strong baselines, covering both traditional retrieval methods and advanced neural reranking models. The baseline configurations slightly differ across datasets.
5.2.1. Baselines on TREC DL 2019
Following (Li et al., 2023a), we include a broad set of competitive baselines for TREC DL 2019, including:
  • Traditional IR: BM25 implemented via Anserini (Yang et al., 2018).
  • Neural IR models: CO-PACRR (Hui et al., 2018), TK (Hofstätter et al., 2020), TKL (Hofstätter et al., 2020), PARADE (Li et al., 2020), RoBERTa-based aggregators (FirstP, MaxP) (Liu et al., 2019), and sparse attention models such as Sparse-Transformer (Child et al., 2019), Longformer-QA (Beltagy et al., 2020), Transformer-XH (Zhao et al., 2020), and QDS-Transformer (Jiang et al., 2020).
  • Previous SOTA reranking system: idst_bert_r1 (Craswell et al., 2020).
  • Previous key block selection models: IDCM (Hofstätter et al., 2021) and KeyB (Li et al., 2023a).
  • LLM-based reranker: RankLLaMA (Ma et al., 2024), along with our implemented variants RankLLaMA-MaxP and RankLLaMA-AvgP for segment-level score aggregation.
5.2.2. Baselines on TREC DL 2023, Robust04, and MLDR-zh
For these datasets, we compare the following unified set of baselines:
  • BM25: implemented with Anserini for English datasets, and adapted with word segmentation for MLDR-zh.
  • KeyB models: KeyB with BM25-based and BinB-based block selection, using either BERT or Chinese-BERT encoders depending on language.
  • LLM-based reranker: RankLLaMA, RankLLaMA-MaxP, and RankLLaMA-AvgP.
5.2.3. RankLLaMA-MaxP and RankLLaMA-AvgP
To enable a similar comparison with the block-based design of KeyB2, we implement two extended variants of RankLLaMA that adopt the same document segmentation strategy as KeyB2. Each document is first divided into same blocks as KeyB2. The final document-level relevance score is obtained by aggregating the block-level scores: RankLLaMA-MaxP selects the highest block score, while RankLLaMA-AvgP averages all block scores. These two aggregation variants serve to contrast the simple score aggregation schemes against the select-then-combine-then-evaluate paradigm employed by KeyB2.
@@ 5.3. Experimental Design and Implementation Details
5.3.1. Experimental Design
To evaluate the proposed KeyB2 approach, we conduct experiments on multiple datasets with both fine-tuning scenarios and zero-shot transfer. All models except BM25 are experimented the reranking effectiveness. We first fine-tune the models on the TREC 2019 DL document ranking dataset and compare their performance with various baselines. To ensure fair comparison with RankLLaMA, which utilizes the Llama2777https://huggingface.co/meta-llama/Llama-2-7b-hf 7B model, we initially fine-tune KeyB2 using the same Llama2 backbone, denoted as KeyB2(Llama2). To further investigate scalability, we additionally fine-tune KeyB2 on the Llama3888https://huggingface.co/meta-llama/Meta-Llama-3-8B 8B model, denoted as KeyB2(Llama3).
The KeyB2 framework supports three block selection strategies: BM25, bi-encoder, and cross-encoder. We train KeyB2 under all three selectors, resulting in variants such as KeyB2(Llama2)BM25, KeyB2(Llama2)bi, and KeyB2(Llama2)cross. We further conduct additional fine-tuning experiments on the TREC 2023 DL dataset to test transferability across MS MARCO v1 and v2 collections.
For zero-shot evaluation, we directly apply the KeyB2 models trained on TREC 2019 DL to Robust04 without any additional fine-tuning. Finally, both KeyB2 and baseline models are fine-tuned on the Robust04 and MLDR-zh datasets to assess in-domain ranking performance. For all datasets, BM25 directly produces initial retrieval rankings, while all other models are evaluated as rerankers operating on candidate document lists (Yang et al., 2018).
5.3.2. Implementation Details
All models are implemented in PyTorch 2.4.0 with CUDA 11.8. All training and testing experiments are performed on a NVIDIA A100-40GB GPU unless otherwise noted.
  1. (1)
    Training Data:
    • TREC 2019 DL: 200K training triplets (query, positive, negative) sampled from msmarco-doctrain-top100999https://microsoft.github.io/msmarco/TREC-Deep-Learning-2019.html, based on MS MARCO v1. The qrel file is used to identify positives, together with the top100 file to sample negatives from the top-100 retrieved documents, forming the training triplets.
    • TREC 2023 DL: 200K triplets constructed from docv2_train_top100.txt.gz [^10] https://microsoft.github.io/msmarco/TREC-Deep-Learning.html, derived from MS MARCO v2. Triplet generation follows the same procedure as TREC 2019 DL.
    • Robust04: 5-fold cross-validation is conducted, with each fold containing 40K triplets. Fold splits follow the CEDR data protocol [^11] https://github.com/Georgetown-IR-Lab/cedr/tree/master/data/robust. For each fold, positives and negatives are formed using the train.pairs and qrel file.
    • MLDR-zh: 10,000 training queries, each paired with one labeled positive and one labeled negative document, as provided by the official dataset annotations.
  2. (2)
    Training Configurations:
    • For KeyB2 models, we apply LoRA (Hu et al., 2021) to efficiently fine-tune the large language models while freezing the backbone weights. LoRA parameters are set to $r=32$ and $\alpha=64$, with a batch size of 2 and gradient accumulation steps of 8, yielding an effective batch size of 16. The learning rate is fixed at $5\times 10^{-5}$, with pairwise hinge loss (Li et al., 2023a), using AdamW optimizer with WarmupDecayLR scheduler, and FP16 mixed-precision training is applied throughout.
    • For RankLLaMA, due to the longer input length (4096 tokens), the batch size is set to 1. We initially attempted fine-tuning RankLLaMA on an NVIDIA H20 GPU (96GB), but it encountered out-of-memory (OOM) errors. Eventually, RankLLaMA models were trained on an A100 using gradient checkpointing to fit into memory. All other hyperparameters are identical to KeyB2.
    • For RankLLaMA-MaxP and AvgP, firstly same input tokens are used as RankLLaMA, then the tokens are segmented into same blocks are KeyB2 models. During training, max or average block score is used to represent document relevant score. All other hyperparameters are identical to KeyB2.
    • For all KeyB(BERT) models, we follow the same training configuration as in (Li et al., 2023a), using a learning rate of $5\times 10^{-5}$, batch size of 4, and gradient accumulation steps of 8, with pairwise hinge loss (Li et al., 2023a). For English benchmarks (e.g., TREC DL19), we initialize from the bert-base-uncased model [^12] https://huggingface.co/google-bert/bert-base-uncased, and for the Chinese benchmark MLDR-zh, we use bert-base-chinese [^13] https://huggingface.co/google-bert/bert-base-chinese as the starting point.
    • All models are trained with the same training data (triplets) for 1 epoch.
@@ 5.3.2. Implementation Details
    • Queries are truncated to a maximum of 32 tokens for all models.
    • For KeyB and KeyB2, block selection reduces documents to 480 tokens, which reduce the input length to approximate 512 tokens.
    • RankLLaMA directly consumes long input up to 4096 tokens, which means approximate 4096-32=4064 document tokens.
@@
    Block Selection Pretrained Models: We utilize language-specific cross-encoders and bi-encoders as block selectors in KeyB2. The detailed model configurations and sources are summarized in Table 3.
@@
    • For English datasets, BM25 is implemented using the Anserini tool (Yang et al., 2018) with default settings ($k_{1}=0.9$, $b=0.4$). For KeyB2 with BM25 selector, we computed the IDF dictionary using scikit-learn on the TREC 2019 DL dataset, with parameters matching Anserini’s defaults.
    • For Chinese dataset MLDR-zh, BM25 baseline is also implemented with Anserini with deault settings. For KeyB2 with BM25 selector, as mentioned in Section 4.3.1, the sentences are separated into Chinese words. Other settings are the same as English datasets.
@@
    • TREC 2019 DL: NDCG@10 and MAP, following prior works (Li et al., 2020, 2023a).
    • TREC 2023 DL is the same as TREC 2019 DL.
    • Robust04: We report P@10, P@20, NDCG@10, NDCG@20, and MAP to evaluate both precision and ranking depth.
    • MLDR-zh: As each query has one positive document within 8 candidates, we report P@1 (first-rank accuracy), MAP (global ranking quality), and NDCG@8 (full list ranking consistency).
    • All models report the reranking results with the metrics. Except Robust04, they rerank the official first-stage results, and on Robust04, they rerank the BM25 top 100 results from Anserini.
  7. (7)
    Efficiency Measurement:
    • Inference latency: Average reranking time for 100 candidate documents per query on a single A100 GPU.
    • GPU memory consumption: Peak memory usage measured during fine-tuning (batch size 1 for both KeyB2 and RankLLaMA).
    • Speed-effectiveness tradeoff: All models are compared based on their Pareto frontier of accuracy versus computational cost.
@@
TREC 2019, 2023 DL & Robust04 MLDR-zh
@@
6. Experimental Results
6.1. Dataset–wise Effectiveness Results
Table 4. Experiment on TREC 2019 DL document reranking benchmark and comparison with sparse attention models and previous select blocks models. Results of KeyB(BERT), LLM part and KeyB2 models are newly trained in this study and other baseline results are taken from (Li et al., 2023a). Best results are in bold.
TREC 2019 Deep Learning Track Document Reranking
@@ 6.1. Dataset–wise Effectiveness Results
RankLLaMA 0.701 0.288
RankLLaMA-MaxP 0.643 0.269
RankLLaMA-AvgP 0.654 0.264
@@
KeyB2(Llama2)BM25 0.713 0.293B
KeyB2(Llama2)cross 0.730RB 0.302B
KeyB2(Llama2)bi 0.710 0.293B
KeyB2(Llama3)BM25 0.718 0.293B
KeyB2(Llama3)cross 0.731RB 0.298B
KeyB2(Llama3)bi 0.711 0.296B
Table 5. Experiment on TREC 2023 DL document reranking benchmark and comparison with sparse attention models and previous select blocks models. Best results are in bold.
TREC 2023 Deep Learning Track Document Reranking
Model NDCG@10 MAP
Baseline models
BM25 0.295 0.105
KeyB(BERT)BM25 0.352 0.123
KeyB(BERT)BinB 0.364 0.128
LLM
RankLLaMA 0.386 0.133
RankLLaMA-MaxP 0.341 0.112
RankLLaMA-AvgP 0.349 0.117
KeyB2 models
KeyB2(Llama2)BM25 0.398B 0.136
KeyB2(Llama2)cross 0.421RB 0.143RB
KeyB2(Llama2)bi 0.411RB 0.141RB
KeyB2(Llama3)BM25 0.403B 0.134
KeyB2(Llama3)cross 0.412B 0.138B
KeyB2(Llama3)bi 0.414RB 0.139B
Table 6. Results on Robust04. Left block: zero-shot transfer; right block: fine-tuned. Best scores in each block are bold.
Zero-shot Fine-tuned
Model P@10 P@20 MAP NDCG@10 NDCG@20 P@10 P@20 MAP NDCG@10 NDCG@20
Baseline models
BM25 0.438 0.363 0.233 0.449 0.424 0.438 0.363 0.233 0.449 0.424
KeyB(BERT)BM25 0.437 0.372 0.223 0.446 0.426 0.489 0.416 0.247 0.504 0.481
KeyB(BERT)BinB 0.470 0.396 0.244 0.485 0.461 0.501 0.423 0.256 0.516 0.490
LLM
RankLLaMA 0.459 0.386 0.229 0.474 0.450 0.468 0.405 0.230 0.476 0.457
RankLLaMA-MaxP 0.426 0.357 0.212 0.436 0.412 0.455 0.384 0.226 0.475 0.447
RankLLaMA-AvgP 0.416 0.362 0.212 0.427 0.413 0.446 0.380 0.223 0.462 0.440
KeyB2 models
KeyB2(Llama2)BM25 0.488R 0.407R 0.240R 0.498R 0.471R 0.527R‡ 0.445RB 0.267R‡ 0.542R‡ 0.514RB
KeyB2(Llama2)cross 0.474 0.402R 0.241R 0.493R 0.469R 0.502R 0.433R 0.255R 0.520R 0.499R
KeyB2(Llama2)bi 0.472 0.399R 0.238R 0.490R 0.466R 0.535RB‡ 0.448RB‡ 0.269R‡ 0.546RB‡ 0.517RB‡
KeyB2(Llama3)BM25 0.489R 0.413RB 0.247R† 0.501R 0.476R 0.564RB 0.475RB 0.282RB 0.576RB 0.547RB
KeyB2(Llama3)cross 0.500RB∗ 0.419RB∗ 0.252R†∗ 0.516RB†∗ 0.489RB†∗ 0.581RB 0.480RB 0.292RB† 0.598RB†∗ 0.559RB
KeyB2(Llama3)bi 0.478R 0.407R 0.244R 0.490 0.469R 0.567RB 0.476RB 0.288RB 0.575RB 0.546RB
Table 7. Results on MLDR Chinese dataset. Best results are in bold.
Model P@1 MAP NDCG@8
Baseline models
BM25 0.201 0.259 0.277
BERT based models
KeyB(BERT)BM25 0.423 0.586 0.685
KeyB(BERT)BinB 0.804R 0.876R 0.907R
LLM
RankLLaMA 0.649 0.755 0.814
RankLLaMA-MaxP 0.374 0.554 0.661
RankLLaMA-AvgP 0.228 0.433 0.567
KeyB2 models
KeyB2(Llama2)BM25 0.864RB 0.910RB 0.932RB
KeyB2(Llama2)cross 0.929RB 0.955RB 0.966RB
KeyB2(Llama2)bi 0.948RB†‡ 0.966RB†‡ 0.974RB†‡
KeyB2(Llama3)BM25 0.909RB 0.942RB 0.956RB
KeyB2(Llama3)cross 0.945RB† 0.966RB† 0.975RB†
KeyB2(Llama3)bi 0.946RB† 0.965RB† 0.974RB†
Significant test: differences are verified with a paired t-test ($p$-value 0.05$\leq$ 0.05≤ 0.05) is used for measuring significance in each table. Superscripts used in all tables for KeyB2 model are: R = significantly better than RankLLaMA; B = significantly better than KeyB(BERT)BinB; / / = significantly better than the BM25 / bi-encoder / cross-encoder selector, respectively. Then we will show the results on the datasets.
Tables 47 list detailed scores for every baseline and all KeyB2 variants on five evaluation settings. Below we summarise, without yet interpreting the research questions, the main numerical facts for each dataset. All relative improvements are reported as percentages “%”, computed against the strongest baseline on the same dataset.
  • TREC DL19: Table 4 shows that KeyB2(Llama2)cross attains 0.730 NDCG@10 and 0.302 MAP. Compared with previous SOTA reranking system idst_bert_r1 (Craswell et al., 2020) with NDCG@10 0.719 and MAP 0.292, the two KeyB2 with cross-encoder models consistently outperform it for both metrics, making KeyB2 models reach new SOTA results on TREC 2019 DL document reranking task. Compared with RankLLaMA (0.701 / 0.288), this is a +4.1 % increase in NDCG@10 and +4.9 % in MAP, achieved with only 512 input tokens instead of 4096. All three selectors with Llama3 exceed both RankLLaMA (including MaxP, AvgP) and the previous best KeyB(BERT)BinB.
  • TREC DL23: On the MS MARCO v2 corpus (Table 5), KeyB2(Llama2)cross yields 0.421 NDCG@10 and 0.143 MAP, corresponding to +9.1 % and +7.5 % over RankLLaMA (0.386 / 0.133). Every selector with Llama2 beats RankLLaMA (including MaxP, AvgP) and KeyB(BERT), confirming robustness under corpus shift.
  • Robust04 (Zero-shot): When models trained on DL19 are evaluated on Robust04 without tuning (Table 6), KeyB2(Llama3)cross achieves 0.516 NDCG@10 and 0.252 MAP. Relative to RankLLaMA (0.474 / 0.229), the gains are +8.9 % NDCG@10 and +10.0 % MAP. Even the BM25 selector surpasses all pooling baselines (RankLLaMA-MaxP / AvgP).
  • Robust04 (Fine-tuned): After five-fold in-domain fine-tuning (Table 6), KeyB2(Llama3)cross reaches 0.598 NDCG@10 and 0.292 MAP. This is +15.9 % and +14.1 % relative to the best non-LLM baseline KeyB(BERT)BinB (0.516 / 0.256), and +25.6 % NDCG@10 score over the RankLLaMA score.
  • MLDR-zh: On the Chinese long-document benchmark (Table 7), KeyB2(Llama2)bi attains 0.948 P@1, 0.966 MAP, and 0.974 NDCG@8. Against RankLLaMA (0.649 / 0.755 / 0.814) this corresponds to +46.1 % P@1, +27.9 % MAP, and +19.7 % NDCG@8. All KeyB2 variants dramatically outperform both LLM (including the MaxP and AvgP models) and KeyB baselines, showing that the framework generalises beyond English corpora.
Across all five evaluation settings, KeyB2 consistently yields sizeable relative improvements over the strongest prior methods, irrespective of language or domain shift. The following subsections (6.2, 6.3 and 6.4) analyse these numbers in depth, addressing RQ2 (effectiveness), RQ4 (efficiency), and RQ3 (selector trade-offs), respectively.
6.2. Cross-Dataset Effectiveness Summary (Answer to RQ2)
Table 8 collates—for each collection—the strongest KeyB2 configuration and the two baselines of RQ2: full-document RankLLaMA and previous best block pre-ranking approach KeyB(BERT)BinB. We report NDCG@10 or NDCG@8 as the main metric, together with relative improvements and significance markers from the detailed result tables.
Table 8. Best KeyB2 variant vs. RankLLaMA and KeyB(BERT)BinB. Gains are relative (%).
Dataset Metric RankLLaMA KeyB(BERT)BinB  KeyB2 (best)
DL19 NDCG@10 0.701 0.697 0.731R,B (+4.3 / +4.9)
DL23 NDCG@10 0.386 0.364 0.421R,B (+9.1 / +15.7)
Robust04 (ZS) NDCG@10 0.474 0.485 0.516R,B (+8.9 / +6.4)
Robust04 (FT) NDCG@10 0.476 0.516 0.598R,B (+25.4 / +15.9)
MLDR-zh NDCG@8 0.814 0.907 0.975R,B (+19.8 / +7.5)